FEL Upgrade Project Weekly Brief July 15-19, 2002

 

 

Highlights:

Another very busy week for the FEL team: On Monday we participated in a briefing to the Navys Vice Chief of Operations; from Monday to Wednesday we were involved in DOEs annual review of Jefferson Labs Science and Technology programs; our first dipole magnets for the Upgrade (GWs and GVs) are being readied for shipment this week; and the optics team put a first assembly together of the outcoupler system for the optical cavity for testing.

We are pleased to congratulate Wael Ibrahim who completed his requirements for a Ph.D degree in Electrical Engineering at Old Dominion University on the subject of "Ultrafast electron dynamics in metals". Wael did work using the FEL under the guidance of Michelle Shinn and his thesis advisor, Prof. Hani El Sayed-Ali, a charter member of our Laser Processing Consortium.

Management:

This week the science and technology of the FEL project were reviewed as part of DOEs annual review of the labs S&T programs. The review team gave complimentary remarks on all of the labs programs and as expected provided advice on how to continue to optimize productivity when resources are constrained. With respect to the FEL program the reviewers were "impressed with the accomplishments of the FEL team given the relatively short life of the program" and also complimented the FEL user program start-up, which although very limited in resources, was "productive" with noteworthy accomplishments in nanotube production, ablations studies and protein dynamics. The first significant demonstration of an "energy recovered linac" with IR Demo was deemed very important for future machine development for both the nuclear science and light source communities. Th synergy of the FEL program with the nuclear physics program at JLab is characterized as "excellent" with the cautionary note that potential conflicts should be avoided or minimized by continued vigilance of management.

The Upgrade project monthly report for the month of June 2002 was completed and distributed to the DOE, ONR and AFRL program managers this week.

Funding ($567k) for the two JTO projects that were funded this year on: (1) injector drive laser development and (2) high power cryogenic optical cavity mirrors was received by our DOE Site Office this week. We thank our contract monitors at JTO ,ONR and DOE for negotiating the bureaucratic path that got the funding to us.

On Monday July 15, Fred Dylla and George Neil participated in a briefing to the Vice Chief of Naval Operations, Admiral Fallon given by Admiral Mathis of NAVSEA and Admiral Cohen of ONR concerning the status and plans for developing FELs for Naval applications. A follow-up briefing to the CNO is being scheduled for next month.

WBS 3 (Beam Physics):

A "People's Injector Subcommittee" (Benson, Hernandez, Douglas) of the commissioning politburo met to discuss injector recovery. Most parameters are well characterized using Demo procedures; a PARMELA study is planned to develop means to lock in the few remaining indeterminants (buncher phase & gradient, SRF cavity 4 phase).

WBS 4 (Injector):

The tapered shielded bellows and shielded / BPM bellows drawings were signed off and submitted to the shop for manufacturing. The bellows and flanges for both are on hand. The cathode support tube is having arc marks polished out in preparation for implantation. Testing continued with W&M on the implantation of a test version of the cathode support tube. Arcing occurred again in the second implantation run. The voltage set point for arcing has been identified and a new test will be performed next week at lower voltages. The mounting supports for the tube will be cleaned from SiO2 before each run.

Gun HVPS - The mounting holes for the HVPS Tank must be enlarged to enable it to be aligned to the Gun Tank. This work is planned for this weekend. The 15 psig relief valve was received. A prototype conditioning resistor was made this week, but it needs additional firing to increase the resistance to 160 Mohm.

WBS 6 (RF):

Quarter HVPS - The hardware, trip limit settings, and PLC code are ready for testing without the 480 VAC connected.

Zone 4 - The cryomodule needs to be retrofitted with a larger valve and several spool pieces to make it identical to zone 2. SRF is working to have all work completed prior to the U-tube party on 7/24/02. Alignment is to be completed on 7/19/02. The network hardware for all RF zones was installed this week. Software is presently trying to get it to operate. The PSS system has not yet been installed for FL04.

WBS 8 (Instrumentation):

The 5 mm hole were successfully placed in the 3 linac beam viewer foils. There are 4 additional foils ready for mounting. The first 4 DV trim magnets were mounted to the girders and are having their input leads configured. The major installation activity for the week was the reworking of the video system. The video crosspoint is being upgraded to 192 video inputs to 16 outputs (these are blocks of 64X16). The field wiring to each of the beam viewer video is included in this effort. The VME video frame grabber software was upgraded, as soon as the video is back up (next week) the viewer calibration files will begin to be created.

We received the F0151 BLM front PCB layout (component placement) and routing has begun. This effort should take a week then we will go out for the four required boards (total 48 channels). The Drive Laser Pulse Controller (DLPC) design review was held. There are a couple of 'features' that need to be worked out, Steve Benson has the action item to work out with G. Neil and M. Shinn. The marching orders are now complete for the entire Machine Protection System.

We have begun an EPICS on Linux effort with 2 for the students. One of the applications is to use a PC/104 board as an IOC and put it in the tunnel to control the IR optical cavity motors and TC guages, the other is to do video processing of the beam viewer signals to have an automated emittance measurement.

WBS 9 (Transport):

Dipoles

Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland CA. The first GW dipole should ship today. (See picture)

The glue down of the shims on the second set of cores is reported to be perfect, variation at levels less than .001.

No word on the 6 cores for the second batch of 6 Magnets for UV.

Injector Dipoles (DU/DV)

The two GU cores were received by Wang NMR and are in trial assembly.

At Wang NMR they are waiting for the delivery of the missing electrical connectors from a local supplier in order to complete the assembly of the GVs. They now predict shipping to us next week.

Arc 180 Degree Dipoles (GY)

At Wang NMR, the fourth GY Coil is starting to be wound.

GY Coil Potting fixture remains in fabrication.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

Process Equipment Co. (PECo) has checked the gap of the three remaining GQ cores and is doing the final machining on them to get the tooling balls in the right locations.

The shim pieces for the energy change were are ground to size.

At Wang NMR, the GX coil winding continues. Potting fixtures are in preparation

Quadrupoles

Trim Quad (QT)

We have run the first unit and have found that the side coils are not as well cooled as we would like. We are doing some tests with cooling plate to bring their temperature rise down..

Sextupole (SF)

We continue checking the sextupole drawings from DULY Research. Finishing them will be delayed by lack of anticipated funds.

Octupole (OT)

Work was halted on this magnet due to lack of anticipated funds.

Beam Line and Vacuum

Master Machine continued work on the Arc Vacuum Chambers.

The X Chamber continues fabrication in the Jlab Shop.

Design of the four chambers for the optical chicane region is at the sign-off stage.

Design of the regions around the wiggler continues.

Design of the stands for sextupoles and trim quads is in checking.

WBS 11 (Optics):

Both optical cavity internal assemblies (e.g., gimbal mounting for the mirrors) arrived last Friday afternoon. The outcoupler assembly has been mounted on one of the test & assembly stands and is getting a thorough test. Measurements of the torque required to move the leadscrew show it well (x2) less than the conservative rating given by the motor manufacturer, yet we had a worm gear begin slipping on the stepper motor shaft. The motor/rotary feedthrough vendor thinks this is a fluke, but we are considering more robust techniques to lock it in place. The translation tests conducted last week and this week were the first opportunity to really use the new stepper motor software, and some bugs appeared. These are being worked through. The OCMMS beamline parts were accepted (only one error requires some minor remachining) and one is mounted in the OCMMS test stand. New bids on the optical cavity vacuum vessels will be received on Monday (7/22). Our first attempts to measurement the outgassing rates of the LVDTs were stymied by an inability to even pump below 0.1 milliTorr, so we had to disassemble one to determine the problem. We found parts that had been written with ink, etc, and will start rechecking them after we clean the parts.

The prototype backplane cooled mirror holder and a test mirror are mounted in the Mirror Test Stand with water fittings connected. We anticipate testing to begin late today.

The ultrafast laser system lost the ability to synchronize to an external RF reference. Our troubleshooting determined that the coarse cavity length adjustment motor had failed. The vendor was able to expedite the repair so we can now continue our tests and measurements.